OpenStudy (anonymous):

how do you simplify 2-2xy^2/2x^2y

OpenStudy (anonymous):

Is it: \[2-\frac{ 2xy ^{2} }{ 2x ^{2}y }\]?

OpenStudy (anonymous):

\[\frac{ 2-2xy^2 }{ 2x^2y }\]

OpenStudy (anonymous):

Factor a 2...

OpenStudy (anonymous):

we just separate the numerators \[\frac{ 2-2xy^{2} }{ 2x ^{2}y }\] = \[\frac{ 2 }{ 2x ^{2} y} - \frac{ 2xy ^{2} }{ 2x ^{2}y }\] then we simplify it by general cancellations and division and obtain \[\frac{ 1 }{ x ^{2}y } - \frac{ y }{ x }\]

OpenStudy (anonymous):

it says the answer is 1-xy^2 over x^2y and i dont see how that = the same thing

OpenStudy (anonymous):

\[\frac{ 2-2xy ^{2} }{ 2x ^{2} y}=\frac{ 2(1-xy ^{2}) }{ 2x ^{2}y }=\frac{ 1-xy ^{2} }{ x ^{2}y }\]

OpenStudy (anonymous):

ok i see what i was doing. I factored out the 2 and made it 2(1-1) instead of seeinng that it is 2(1-xy^2) thank you

OpenStudy (anonymous):

No prob